Circular No. 7990

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002gf, 2002gg, AND 2002gh Further to IAUC 7983, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and R. Rudas report the discovery of three apparent supernovae on unfiltered Palomar NEAT CCD images.

SN  2002 UT  α_{2000}  δ_{2000}  Mag.  Offset
2002gf Sept.27.11 20^{h}55^{m}17.^{s}88 − 0^{o}04′21′′.4 19.4 9′′ W, 3′′ N
2002gg Oct. 5.18 22 27 45.93 + 6 55 44.7 20.0 6′′ SE
2002gh Oct. 5.41 3 05 29.46 − 5 21 57.0 19.3 8′′.7 E, 6′′ N

Additional unfiltered magnitudes (the prediscovery limiting magnitudes are given for S/N = 5): SN 2002gf, 2001 Sept. 14, [21.0; 2002 Sept. 4, [21.0; Sept. 28, 19.3. SN 2002gg, 2001 Sept. 25, [21; 2002 Sept. 4, [21; Oct. 6, 20.1. SN 2002gh, 2001 Sept. 27, [21.2; Dec. 14, [21.2; 2002 Oct. 6, 19.2.

SUPERNOVA 2002ge IN NGC 7400 M. Salvo, Australian National University (ANU); and T. Hua, Labora-

toire d’Astrophysique de Marseille, report that two spectra (ranges 460–560 and 620–720 nm) of SN 2002ge (cf. IAUC 7987) were taken with the ANU 2.3-m telescope (+ Double Beam Spectrograph) at Siding Spring Observatory on Oct. 9.52 UT. Although the wavelength ranges are rather small, the two spectra suggest a resemblance to that of a type-Ia supernova ∼ 10 days after maximum light.

B.-G. Andersson, Johns Hopkins University; S. Potter, South African Astronomical Observatory (SAAO); P. Garnavich, University of Notre Dame; and G. Sonneborn, Goddard Space Flight Center, NASA, report that calibrated spectra of SN 2002ge (resolution 0.2 nm), obtained on Oct. 8.94 UT with the SAAO 1.9-m telescope, show an average flux of 10^{−14} erg s^{−1} cm^{−2} Å^{−1}. One spectrum (range 350–550 nm) is that of a type-I supernova. They add: “The flux deficit around 410–420 nm is similar to that of the type-Ic SN 1994I, a few days before maximum, but subluminous type-Ia supernovae can also show this depression.”

SUPERNOVA 2002gd IN NGC 7537 CCD V magnitude estimates for SN 2002gd from A. Klotz (cf. IAUC

7986) and C. Jasinski (J), with the background-galaxy light now subtracted: Sept. 15.95 UT, [18.0; Oct. 5.87, 17.9 (discovery image, revised); 6.83, 17.7; 7.83, 17.3 (J); 8.78, 17.5.

2002 October 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green